• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Peter L

    Log entry time 00:31:16 on July20,2005

    Entry number 148669

    keyword=Swing Shift Summary

    We started down for beam studies and injector work. About 7pm they gave us control of the beam. David and Sharish worked on the compton and were very pleased with the tune so we decided to go ahead with the raster settings. The first two settings were fine as they were but the 7x7 needed to be changed.

    RUN Spot++ name Actual SETTING
    2164 2x2 gamma 3x3 3.8x4.75
    2163 3.5x3.5 beta 5x5 7.5x11.0
    2169 4.5x4.5 alpha 7x7 9.8x13.7

    2156,2157,2158,2159,2560,2561,2562,2165,2166,2167,2168 making adjustments to spot size.

    Next we proceeded with the target boiling studies.
    Run
    3695 7x7(alpha) 35uA final 2000ev at 35uA
    3696 7x7(alpha) 35uA pedestal
    3697 7x7(alpha) 10uA w/ped
    3698 5x5(beta) 10uA w/ped
    3699 5x5(beta) 35uA w/ped
    3700 3x3(gamma) 35uA w/ped
    3701 3x3(gamma) 10uA w/ped

    David took control of the beam to finish Compton studies so we started a profile scan to place the peak.

    3702 rough profile scan
    3703 junk
    3704 detailed profile scan at 25uA

    PROBLEM: we cannot move to Lead Target because the move box refuses to appear. We can move to any other target.